# Theoretical study on $\Xi\alpha$ correlation function

### v1

@@ I Introduction
Understanding the properties of the Hyperon ($Y$)-Nucleon ($N$) interactions is a long standing problem in the hypernuclear physics. Among others, the $\Xi N$ interaction in the strangeness $S=-2$ sector is important to pin down the role of $\Xi$ in the neutron star and possible existence of the $H$ dibaryon [1]. There are four independent componetns in the $s$-wave $\Xi N$ interaction, ${}^{11}S_{0},{}^{31}S_{0},{}^{13}S_{1},{}^{33}S_{1}$ where the notation ${}^{2I+1,2s+1}L_{J}$ is adopted with the spin $s$, isospin $I$, and the total angular momentum $J$.
Through the recent experimental observations of the $\Xi$ hypernuclei [2, 3, 4], the $\Xi N$ interaction is considered to be moderately attractive in total. This is consistent with the recent theoretical analysis with the chiral effective field theories [5, 6] and the lattice QCD calculations [7, 8]. On the other hand, the scattered results of the theoretical predication on the light $\Xi$ hypernuclei [9, 10, 11, 12, 13, 14] indicate that the detailed understanding of the $\Xi N$ interactions is still insufficient and further constraints through the experimental observables are required.
@@
For the study of the $\Xi N$ interactions, $p\Xi^{-}$ correlation functions from the $pp$ [20] and $p$Pb collisions [19] are measured by the ALICE collaboration. The corresponding theoretical studies can be found in Refs. [26, 27]. In Ref. [27], the ALICE data are shown to be consistent with the theoretical correlation function calculated with the lattice QCD $\Xi N$-$\Lambda\Lambda$ coupled-channel potential [8]. The measured correlation function is also in good agreement with the result in Ref. [26], which is evaluated with the chiral effective field theory [5], although the spin-isospin components of the $\Xi N$ interaction exhibit different properties from the lattice QCD potential. This is because the enhancement of the $\Xi N$ correlation is originated mainly in the strong attraction in the ${}^{11}S_{0}$ component, which shows similar behavior in both potentials. Thus, with these studies, strongly attractive nature of the ${}^{11}S_{0}$ interaction is well confirmed, while the other components still lack the thorough understanding.
@@ Acknowledgements.
We thank Emiko Hiyama for providing the $\Xi\alpha$ folding potential and for the helpful support. We also thank Kouichi Hagino, Yudai Ichikawa, Johann Haidenbauer, Andreas Nogga, and Hoai Le for useful discussions and comments. This work was supported in part by the Grants-in-Aid for Scientific Research from JSPSs (Grants No. JP23H05439, and No. JP22K03637), by JST, the establishment of university fellowships towards the creation of science technology innovation, Grant No. JPMJFS2123, by JST SPRING, Grant Number JPMJSP2110, and by the Deutsche Forschungsgemeinschaft (DFG) and the National Natural Science Foundation of China (NSFC) through the funds provided to the Sino-German Collaborative Research Center “Symmetries and the Emergence of Structure in QCD” (NSFC Grant No. 12070131001, DFG Project-ID 196253076 – TRR 110).

### v2

@@ I Introduction
Understanding the properties of the Hyperon ($Y$)-Nucleon ($N$) interactions is a long standing problem in the hypernuclear physics. Among others, the $\Xi N$ interaction in the strangeness $S=-2$ sector is important to pin down the role of $\Xi$ in the neutron star and possible existence of the $H$ dibaryon [1]. There are four independent components in the $s$-wave $\Xi N$ interaction, ${}^{11}S_{0},{}^{31}S_{0},{}^{13}S_{1},{}^{33}S_{1}$ where the notation ${}^{2I+1,2s+1}L_{J}$ is adopted with the spin $s$, isospin $I$, and the total angular momentum $J$.
Through the recent experimental observations of the $\Xi$ hypernuclei [2, 3, 4], the $\Xi N$ interaction is considered to be moderately attractive in total. This is consistent with the recent theoretical analysis with the chiral effective field theories [5, 6, 7] and the lattice QCD calculations [8, 9]. On the other hand, the scattered results of the theoretical predication on the light $\Xi$ hypernuclei [10, 11, 12, 13, 14, 15, 16, 17] indicate that the detailed understanding of the $\Xi N$ interactions is still insufficient and further constraints through the experimental observables are required.
@@
For the study of the $\Xi N$ interactions, $p\Xi^{-}$ correlation functions from the $pp$ [23] and $p$Pb collisions [22] are measured by the ALICE collaboration. The corresponding theoretical studies can be found in Refs. [29, 30, 31, 32]. In Ref. [30], the ALICE data are shown to be consistent with the theoretical correlation function calculated with the lattice QCD $\Xi N$-$\Lambda\Lambda$ coupled-channel potential [9]. The measured correlation function is also in good agreement with the results in Refs. [29, 31, 32], which are evaluated with the chiral effective field theory [5, 7], although the spin-isospin components of the $\Xi N$ interaction exhibit different properties from the lattice QCD potential. This is because the enhancement of the $\Xi N$ correlation is originated mainly in the strong attraction in the ${}^{11}S_{0}$ component, which shows similar behavior in both potentials. Thus, with these studies, strongly attractive nature of the ${}^{11}S_{0}$ interaction is well confirmed, while the other components still lack the thorough understanding.
@@ Acknowledgements.
We thank Emiko Hiyama for providing the $\Xi\alpha$ folding potential and for the helpful support. We also thank Kouichi Hagino, Yudai Ichikawa, Johann Haidenbauer, Andreas Nogga, Hoai Le, and Avraham Gal for useful discussions and comments. This work was supported in part by the Grants-in-Aid for Scientific Research from JSPSs (Grants No. JP23H05439, and No. JP22K03637), by JST, the establishment of university fellowships towards the creation of science technology innovation, Grant No. JPMJFS2123, by JST SPRING, Grant Number JPMJSP2110, and by the Deutsche Forschungsgemeinschaft (DFG) and the National Natural Science Foundation of China (NSFC) through the funds provided to the Sino-German Collaborative Research Center “Symmetries and the Emergence of Structure in QCD” (NSFC Grant No. 12070131001, DFG Project-ID 196253076 – TRR 110).